a rectangular prism has a volume of 3x^2+8x^2-45x-50. if the height is (x+5) find its length and width
Answer by ewatrrr(24785)

 
Hi
Using synthetic Division
-5 3 8 -45 -50
-15 35 50
3 -7 -10 0
3x^2 - 7x - 10 = 0
(3x -10)(x + 1) = 0
L and width are: (3x-10) and (x+1)